OpenStudy (swaugh):

What is the distance from negative three comma one to negative one comma five Round your answer to the nearest hundredth.

OpenStudy (anonymous):

You would have to use the distance formula, when you plug in those coordinates it would equal to radical 20

OpenStudy (campbell_st):

if you plot the points join them the form a right triangle  you can use pythagoras' theorem to find the distance hope it helps

OpenStudy (swaugh):

i dont know how to do pythagoras therem

OpenStudy (campbell_st):

well if you don't know pythagoras then it makes it hard to find the distance  Pythagoras' theorem can be used to find the longest side in a right triangle the theorem is \[a^2 + b^2 = c^2\] so after you calculate the value \[c = \sqrt{a^2 + b^2}\] so find the distances between x = -3 and x = -1 let this be a and find the distance y = 1 to y = 5 and let this be b now you can find the distance between the 2 points